I'm not sure how to calculate the breakeven

Assuming monetary benefits of an information system at $85,000 per year, one-time costs of $75,000, recurring costs of $35,000 per year, a discount rate of 12 percent,and a five-year time horizon, calculate the net present value of these costs and benefits of an information system. Also calculate the overall return on investment ofthe project and then present a break-even analysis. At what point does break-even occur?
Answer #1
Monetary benefits system= $85000
One Time costs = $75000
Recurring cost= $35000
Discount rate= 12 percent------Cash flow and discount per year is calculated as

to-----75000

t1 ---(85000-35000) / 1.12 = 44,642
t2---- (85000-35000) / 1.12^2 = 39,859
t3 ---(85000-35000) / 1.12^3 = 35,589
t4--- (85000-35000) / 1.12^4 = 31,775
t5---- (85000-35000) / 1.12^5 = 28,371

------ NPV =(t1+t2+t3+t4+t5)-$75000 = $105,236

So, breaks-even point is after 2 years. (t1+t2) =$84,501 nearly equalto monetary benefits
Here, the NPV at that point is positive, whenthe costs and benefits were throughout the year,

answered by: swischeese